a train travelling at 20km/hr is approaching a platform a bird is siiting on a pole on the platform when the train is at a distance of 2 km from the pole brakes are applied which produce a uniform deceleration in it at that instant the bird flies towards the train at 60km/hr and after touching the nearest point on the train flies back to the pole and then flies towards the train and continues repeating itself calculate how much distance will the bird have flown before the train stopps
 

Grade:10

1 Answers

Arun
6 years ago
Dear Kanchan
 

Initial velocity of train = 20km/hr. Final velocity of the train = 0 km/hr as it comes to rest finally.

distance =2 km

 

V² = u² +2as

By putting the values

0 = 400 +4a

a = -100km/hr²

Now to calculate time taken by train to travel 2km to reach the platform.

V = u +at

0 = 20 - 100t

t = 1/5 hr

Distance = speed* time = 60* (1/5) = 12 km

 

Regards

Arun (askIITians forum expert)